# SeaS: Few-shot Industrial Anomaly Image Generation with Separation and Sharing Fine-tuning

### v1

@@ 3.3 Refined Mask Prediction
High-fidelity pixel-wise annotations of anomalies play an important role in boosting segmentation models. However, existing methods, such as DFMGAN (Duan et al., 2023) and AnomlayDiffusion (Hu et al., 2024), produce anomaly masks that are not tightly matched with generated anomalies, which is insufficient for training segmentation model. To address this issue, we design a cascaded Refined Mask Prediction (RMP) branch, which is grafted onto the U-Net trained according to SeaS (mentioned in Sec. 3.2). As shown in Fig. 4, RMP consists of two steps, firstly capturing discriminative features from U-Net and secondly combining it with high-resolution features of VAE decoder to generate anomaly-matched masks.
Figure 4: The Refined Mask Prediction (RMP) branch during inference. The Coarse Feature Extraction utilizes features from the up-2 and up-3 layers of U-Net Decoder to extract coarse features. The cascaded Mask Refinement Module (MRM) further obtains the mask accurately aligned with the anomaly with the assistance of high-resolution features of the VAE Decoder.
@@
Mask Refinement Module. Directly upsampling the coarse feature $\hat{F}$ to high resolution will result in a loss of anomaly details. Therefore, we design the Mask Refinement Module (MRM) to refine the coarse feature $\hat{F}$ in a progressive manner. As shown in Fig. 4, each MRM takes in two features, i.e., the high-resolution features from VAE and the discriminative feature to be refined. Firstly, the discriminative feature is upsampled to align with the high-resolution features of VAE. To preserve the discriminative ability, the upsampled feature is processed through two chained convolution blocks for capturing multi-scale anomaly features and a $1\times 1$ convolution for capturing local features. These features are then summed, and multiplied with the VAE features element-wisely to enhance the anomalies’ boundary. Finally, MRM employs a $3\times 3$ convolution to fuse the added features and output a refined feature.
@@
Loss Functions. During training, we use $x_{\text{df}}$ and $x_{\text{ob}}$ as inputs. For $x_{\text{df}}$ we obtain the coarse mask $\hat{M}_{\text{df}}\in\mathbb{R}^{64\times 64\times 2}$ from the Coarse Feature Extraction and $\hat{M}^{\prime}_{\text{df}}$ after the MRMs. Similarly, for $x_{\text{ob}}$ we obtain the $\hat{M}_{\text{ob}}\in\mathbb{R}^{64\times 64\times 2}$ from Coarse Feature Extraction and directly upsample it to the original resolution, denoted as $\hat{M}^{\prime}_{\text{ob}}\in\mathbb{R}^{512\times 512\times 2}$. Then we conduct the supervision on both low-resolution and high-resolution predictions as,
$\mathcal{L}_{\mathcal{M}}=\mathcal{F}(\hat{M}_{\text{df}},\mathbf{M}_{\text{df}})+\mathcal{F}(\hat{M}_{\text{ob}},\mathbf{M}_{\text{ob}})+\mathcal{F}(\hat{M}_{\text{df}}^{\prime},\mathbf{M}_{\text{df}}^{\prime})+\mathcal{F}(\hat{M}_{\text{ob}}^{\prime},\mathbf{M}_{\text{ob}}^{\prime})$ (7)
where $\mathcal{F}$ indicates the Focal Loss (Lin et al., 2017). $\mathbf{M}_{\text{ob}}\in\mathbb{R}^{64\times 64\times 1}$ and $\mathbf{M}_{\text{ob}}^{\prime}\in\mathbb{R}^{512\times 512\times 1}$ are used to suppress noise in normal images, with each pixel value set to 0. $\mathbf{M}_{\text{df}}\in\mathbb{R}^{64\times 64\times 1}$ and $\mathbf{M}_{\text{df}}^{\prime}\in\mathbb{R}^{512\times 512\times 1}$ are the ground truth masks of abnormal images. More ablation studies on the effect of normal images in training RMP branch are shown in Tab. 27 and Fig. 16 in appendix A.8.
@@ 3.4 Inference
During the generation of the abnormal image-mask pairs, aiming further to ensure the global consistency of the abnormal image, we random select a normal image $x_{\text{ob}}$ from $X_{\text{ob}}$ as input, and add random noise to $x_{\text{ob}}$, which resulting in an initial noisy latent $\hat{z_{0}}$. Next, $\hat{z_{0}}$ is input into the U-Net for noise prediction, with the process guided by the conditioning vector $\textbf{e}_{\text{df}}$ (mentioned in Eq. 4). In the final three denoising steps, the RMP branch (Sec. 3.3) leverages the features from the U-Net decoder and VAE decoder to generate the final anomaly mask. Specifically, we average the refined anomaly mask from these steps to obtain the refined mask $\hat{M}_{\text{df}}^{\prime}\in\mathbb{R}^{512\times 512\times 2}$. Then we take the threshold $\tau$ for the second channel of $\hat{M}_{\text{df}}^{\prime}$ to segment the final anomaly mask $M_{\text{df}}\in\mathbb{R}^{512\times 512\times 1}$. The effect of $\tau$ on the downstream segmentation models is shown in Tab. 29 in appendix A.8. In the last denoising step, the output of the generation model is used as the generated abnormal image.
Category CDC Crop& Paste SDGAN Defect- GAN DFMGAN Anomaly Diffusion Ours
IS $\uparrow$ IC-L $\uparrow$ IS $\uparrow$ IC-L $\uparrow$ IS $\uparrow$ IC-L $\uparrow$ IS $\uparrow$ IC-L $\uparrow$ IS $\uparrow$ IC-L $\uparrow$ IS $\uparrow$ IC-L $\uparrow$ IS $\uparrow$ IC-L $\uparrow$
bottle 1.52 0.04 1.43 0.04 1.57 0.06 1.39 0.07 1.62 0.12 1.58 0.19 1.78 0.21
cable 1.97 0.19 1.74 0.25 1.89 0.19 1.70 0.22 1.96 0.25 2.13 0.41 2.09 0.42
capsule 1.37 0.06 1.23 0.05 1.49 0.03 1.59 0.04 1.59 0.11 1.59 0.21 1.56 0.26
carpet 1.25 0.03 1.17 0.11 1.18 0.11 1.24 0.12 1.23 0.13 1.16 0.24 1.13 0.25
grid 1.97 0.07 2.00 0.12 1.95 0.10 2.01 0.12 1.97 0.13 2.04 0.44 2.43 0.44
hazelnut 1.97 0.05 1.74 0.21 1.85 0.16 1.87 0.19 1.93 0.24 2.13 0.31 1.87 0.31
leather 1.80 0.07 1.47 0.14 2.04 0.12 2.12 0.14 2.06 0.17 1.94 0.41 2.03 0.40
metal_nut 1.55 0.04 1.56 0.15 1.45 0.28 1.47 0.30 1.49 0.32 1.96 0.30 1.64 0.31
pill 1.56 0.06 1.49 0.11 1.61 0.07 1.61 0.10 1.63 0.16 1.61 0.26 1.62 0.33
screw 1.13 0.11 1.12 0.16 1.17 0.10 1.19 0.12 1.12 0.14 1.28 0.30 1.52 0.31
tile 2.10 0.12 1.83 0.20 2.53 0.21 2.35 0.22 2.39 0.22 2.54 0.55 2.60 0.50
toothbrush 1.63 0.06 1.30 0.08 1.78 0.03 1.85 0.03 1.82 0.18 1.68 0.21 1.96 0.25
transistor 1.61 0.13 1.39 0.15 1.76 0.13 1.47 0.13 1.64 0.25 1.57 0.34 1.51 0.34
wood 2.05 0.03 1.95 0.23 2.12 0.25 2.19 0.29 2.12 0.35 2.33 0.37 2.77 0.46
zipper 1.30 0.05 1.23 0.11 1.25 0.10 1.25 0.10 1.29 0.27 1.39 0.25 1.63 0.30
Average 1.65 0.07 1.51 0.14 1.71 0.13 1.69 0.15 1.72 0.20 1.80 0.32 1.88 0.34
@@ 4.1 Experimental Settings
Implementation Details. We train SeaS by fine-tuning the pre-trained Stable Diffusion v1-4 (Rombach et al., 2022). In AIG experiments, we use 60 normal images and 1/3 abnormal images with their corresponding masks per anomaly type for training. During inference, we generate 1,000 abnormal image-mask pairs for a single anomaly type. More details are given in appendix A.3.
@@
Datasets. We conduct experiments on MVTec AD dataset (Bergmann et al., 2019) and MVTec 3D AD dataset (only RGB images) (Bergmann. et al., 2022). MVTec AD dataset contains 15 product categories, each with up to 8 different anomalies, making it suitable for simulating real-world industrial scenarios. MVTec 3D AD dataset includes 10 product categories, each with up to 4 different anomalies. It contains more challenges, i.e., lighting condition variations, product pose variations. Due to the page limitation, results on MVTec 3D AD dataset are given in appendix A.4.
@@
Evaluation Metrics. For AIG, we leverage 2 metrics: the Inception Score (IS) and the Intra-cluster pairwise LPIPS distance (IC-LPIPS)(Ojha et al., 2021). The scarcity of abnormal images hampers the reliability of FID (Heusel et al., 2017) and KID (Bińkowski et al., 2018), as overfitted model (Duan et al., 2023) achieves higher scores. For pixel-level anomaly segmentation and image-level anomaly detection, we report 3 metrics: Area Under Receiver Operator Characteristic curve (AUROC), Average Precision (AP) and $F_{1}$-score at optimal threshold ($F_{1}$-max). In addition, we report Intersection over Union (IoU) for segmentation.
@@ 4.2 Comparison in Anomaly Image Generation
Comparison Methods. We compare SeaS with the current AG and AIG methods on generation fidelity and diversity, such as CDC (Ojha et al., 2021), Crop&Paste (Lin et al., 2021), SDGAN (Niu et al., 2020), Defect-GAN (Zhang et al., 2021a), DFMGAN (Duan et al., 2023) and AnomalyDiffusion (Hu et al., 2024). Then we use Crop&Paste, DRAEM (Zavrtanik et al., 2021), DFMGAN, AnomalyDiffusion and our method to generate anomaly image-mask pairs. These pairs are used to train BiSeNet V2 (Yu et al., 2021), UPerNet (Xiao et al., 2018) and LFD (Zhou et al., 2024a) respectively. Different from AnomalyDiffusion (Hu et al., 2024), which trains one segmentation model per product, we train a unified segmentation model for all the products. We also compare the segmentation results based on SeaS with the state-of-the-art unsupervised anomaly detection methods, such as RealNet (Zhang et al., 2024) and HVQ-Trans (Lu et al., 2023), in appendix A.5.
@@
Anomaly image generation quality. In Tab. 1, we compare SeaS with some state-of-the-art AG and AIG methods on generation fidelity (IS) and diversity (IC-LPIPS). SeaS achieves 1.88 on IS and 0.34 on IC-LPIPS, which demonstrates that our method generates anomaly images with higher fidelity and diversity. We exhibit the generated anomaly images in Fig. 5, the anomaly images generated by our method have higher fidelity (e.g., hazelnut_crack). Compared with other methods, SeaS can generate images with different types, colors, and shapes of anomalies rather than overfitting to the training images (e.g., wood_color and pill_crack). The masks generated by our method are also precisely aligned with the anomaly regions (e.g., toothbrush_defective). More qualitative and quantitative anomaly image generation results are in appendix A.6.
Figure 5: Visualization of the generation results on MVTec AD. The sub-image in the lower right corner is the generated mask, none means that the method cannot generate masks.
@@
Anomaly segmentation and detection. We generate 1,000 image-mask pairs for each anomaly type, and use the image-mask pairs of all products along with all the training normal images to train the unified segmentation model, rather than training separate segmentation models for each product. We test the models on the rest images of the testing set of MVTec AD, which are not included in the training set for generation. The results are given in Tab. 2. All the methods are trained using the same number of images and the same training settings, detailed in appendix A.7. The segmentation results consistently demonstrate that our method outperforms others across all the segmentation models, with an 11.17% average improvement on IoU. We show the segmentation anomaly maps in Fig. 6. By using our generated image-mask pairs to train BiSeNet V2, there are fewer false positives in wood_combined and fewer false negatives in bottle_contamination and carpet_cut. In addition, we use the maximum value of the segmentation anomaly map as the image-level anomaly score for anomaly detection. We report the image-level metrics in Tab. 3, and our method achieves a 2.77% gain on image-AUROC. More qualitative comparison results on anomaly segmentation are in appendix A.9 and appendix A.10.
@@
Figure 6: Qualitative anomaly segmentation results with BiSeNet V2 on MVTec AD.
Table 2: Comparison on anomaly segmentation on MVTec AD.
Model DRAEM DFMGAN AnomalyDiffusion Ours
AUROC AP $F_{1}$-max IoU AUROC AP $F_{1}$-max IoU AUROC AP $F_{1}$-max IoU AUROC AP $F_{1}$-max IoU
BiSeNet V2 81.37 38.90 42.62 39.39 94.57 60.42 60.54 45.83 96.27 64.50 62.27 42.89 97.21 69.21 66.37 55.28
UPerNet 83.21 42.78 45.97 42.03 92.33 57.01 56.91 46.64 96.87 69.92 66.95 50.80 97.87 74.42 70.70 61.24
LFD 76.41 40.99 43.91 35.61 94.91 67.06 65.09 45.49 96.30 69.77 66.99 45.77 98.09 77.15 72.52 56.47
Average 80.33 40.89 44.17 39.01 93.94 61.50 60.85 45.99 96.48 68.06 65.40 46.49 97.72 73.59 69.86 57.66
Table 3: Comparison on image-level anomaly detection on MVTec AD.
Model DRAEM DFMGAN AnomalyDiffusion Ours
AUROC AP $F_{1}$-max AUROC AP $F_{1}$-max AUROC AP $F_{1}$-max AUROC AP $F_{1}$-max
BiSeNet V2 89.87 93.51 89.97 90.90 94.43 90.33 90.08 94.84 91.84 96.00 98.14 95.43
UPerNet 89.45 93.92 89.66 90.74 94.43 90.37 96.62 98.61 96.21 98.29 99.20 97.34
LFD 87.94 93.41 88.65 91.08 95.40 90.58 95.15 97.78 94.66 95.88 97.89 95.15
Average 89.09 93.61 89.43 90.91 94.75 90.43 93.95 97.08 94.24 96.72 98.41 95.97

### v2

@@ 3.3 Refined Mask Prediction
The design of the separated and shared approach enables U-Net to simultaneously model the different degrees of variations in both normal products and anomalies, representing the discriminative features for mask prediction. To further obtain pixel-accurate masks, we design a cascaded Refined Mask Prediction (RMP) branch, which is grafted onto the U-Net trained within SeaS (mentioned in Sec. 3.2). As shown in Fig. 4, RMP consists of two steps, firstly capturing discriminative features from U-Net and secondly combining them with high-resolution features of VAE decoder to generate anomaly-matched masks.
@@
Mask Refinement Module. Directly upsampling the coarse feature $\hat{F}$ to high resolution will result in a loss of anomaly details. Therefore, we design the Mask Refinement Module (MRM) to refine the coarse feature $\hat{F}$ in a progressive manner. As shown in Fig. 4, each MRM takes in two features, i.e., the high-resolution features from VAE and the discriminative feature to be refined. Firstly, the discriminative feature is upsampled to align with the high-resolution ones of VAE. To preserve the discriminative ability, the upsampled feature is processed through two chained convolution blocks for capturing multi-scale anomaly features and a $1\times 1$ convolution for capturing local features. These features are then summed and multiplied with the VAE features element-wisely to enhance the anomalies’ boundary. Finally, MRM employs a $3\times 3$ convolution to fuse the added features and outputs a refined feature.
@@
Loss Functions. During training, we use $x_{\text{df}}$ and $x_{\text{ob}}$ as inputs. For $x_{\text{df}}$, we obtain the coarse mask $\hat{M}_{\text{df}}\in\mathbb{R}^{64\times 64\times 2}$ from the Coarse Feature Extraction and $\hat{M}^{\prime}_{\text{df}}$ after the MRMs. Similarly, for $x_{\text{ob}}$, we obtain the $\hat{M}_{\text{ob}}\in\mathbb{R}^{64\times 64\times 2}$ from Coarse Feature Extraction and directly upsample it to the original resolution, denoted as $\hat{M}^{\prime}_{\text{ob}}\in\mathbb{R}^{512\times 512\times 2}$. Then we conduct the supervision on both low-resolution and high-resolution predictions as,
$\mathcal{L}_{\mathcal{M}}=$ $\mathcal{F}(\hat{M}_{\text{df}},\mathbf{M}_{\text{df}})+\mathcal{F}(\hat{M}_{\text{ob}},\mathbf{M}_{\text{ob}})+$ (7)
$\mathcal{F}(\hat{M}_{\text{df}}^{\prime},\mathbf{M}_{\text{df}}^{\prime})+\mathcal{F}(\hat{M}_{\text{ob}}^{\prime},\mathbf{M}_{\text{ob}}^{\prime})$
where $\mathcal{F}$ indicates the Focal Loss [24]. $\mathbf{M}_{\text{ob}}\in\mathbb{R}^{64\times 64\times 1}$ and $\mathbf{M}_{\text{ob}}^{\prime}\in\mathbb{R}^{512\times 512\times 1}$ are used to suppress noise in normal images, with each pixel value set to 0. $\mathbf{M}_{\text{df}}\in\mathbb{R}^{64\times 64\times 1}$ and $\mathbf{M}_{\text{df}}^{\prime}\in\mathbb{R}^{512\times 512\times 1}$ are the ground truth masks of abnormal images. More ablation studies on the effect of normal images in training RMP branch are shown in Tab. 16 and Fig. 11 in appendix A.5.
@@ 3.4 Inference
During the generation, aiming further to ensure the global consistency of the normal products, we random select a normal image $x_{\text{ob}}$ from $X_{\text{ob}}$ as input, and add random noise to $x_{\text{ob}}$, which resulting in an initial noisy latent $\hat{z_{0}}$. Next, for the generation of abnormal images, $\hat{z_{0}}$ is input into the U-Net for noise prediction, with the process guided by the conditioning vector $\textbf{e}_{\text{df}}$ (mentioned in Eq. 4), which is corresponding to the whole UA Text Prompt $\mathcal{P}$. For generating normal images to further enhance unsupervised AD methods, we use the conditioning vector $\textbf{e}_{\text{ob}}$ corresponding to the normal tokens of $\mathcal{P}$. Regarding the masks corresponding to anomalies, in the final three denoising steps, the RMP branch (Sec. 3.3) leverages the features from the U-Net decoder and VAE decoder to generate the final anomaly mask. Specifically, we average the refined anomaly mask from these steps to obtain the refined mask $\hat{M}_{\text{df}}^{\prime}\in\mathbb{R}^{512\times 512\times 2}$. Then we take the threshold $\tau$ for the second channel of $\hat{M}_{\text{df}}^{\prime}$ to segment the final anomaly mask $M_{\text{df}}\in\mathbb{R}^{512\times 512\times 1}$. The effect of $\tau$ on the downstream supervised segmentation models is shown in Tab. 18 in appendix A.5. In the last denoising step, the output of the generation model is used as the generated abnormal image.
@@ 4 Experiments
Table 1: Comparison on IS and IC-LPIPS on MVTec AD, VisA, and MVTec AD 3D. Bold indicates the best performance.
Methods MVTec AD VisA MVTec 3D AD
[…]
Crop&Paste[23] 1.51 0.14 - - - - - - - - - -
SDGAN[28] 1.71 0.13 - - - - - - - - - -
Defect-GAN[42] 1.69 0.15 - - - - - - - - - -
DFMGAN[11] 1.72 0.20 0.12 0.14 1.25 0.25 0.24 0.05 1.80 0.29 0.19 0.08
AnomalyDiffusion[17] 1.80 0.32 - 0.12 1.26 0.25 - 0.04 1.61 0.22 - 0.07
SeaS 1.88 0.34 0.04 0.18 1.27 0.26 0.02 0.06 1.95 0.30 0.06 0.09
@@ 4.1 Experimental Settings
Implementation Details. We train SeaS by fine-tuning the pre-trained Stable Diffusion v1-4 [31]. In anomaly image generation experiments, we use $60$ normal images and $\frac{1}{3}$ masked anomaly images for each anomaly type in training. We train one generative model per product, covering all anomaly types. During inference, we generate $1,000$ anomaly image-mask pairs for a single anomaly type. More details are given in appendix A.3.
Figure 5: Visualization of the generation results on MVTec AD, VisA and MVTec 3D AD. The sub-image in the lower right corner is the generated mask, none means that the method cannot generate masks.
Figure 6: Visualization of the generated normal images on MVTec AD, VisA and MVTec 3D AD.
@@
Datasets. We conduct experiments on MVTec AD dataset [3], VisA dataset[47], and MVTec 3D AD dataset (only RGB images) [4]. MVTec AD dataset contains $15$ product categories, each with up to $8$ different anomalies. VisA dataset covers $12$ objects in $3$ domains. MVTec 3D AD dataset includes $10$ product categories, each with up to $4$ different anomalies. It contains more challenges, i.e., lighting condition variations, and product pose variations.
Table 2: Comparison on combining generated anomalies with synthesis-based anomaly detection methods across multiple datasets.
Segmentation Models MVTec AD VisA MVTec 3D AD
Image-level Pixel-level Image-level Pixel-level Image-level Pixel-level
AUROC AP $F_{1}$-max AUROC AP $F_{1}$-max IoU AUROC AP $F_{1}$-max AUROC AP $F_{1}$-max IoU AUROC AP $F_{1}$-max AUROC AP $F_{1}$-max IoU
DRAEM [41] 98.00 98.45 96.34 97.90 67.89 66.04 60.30 86.28 85.30 81.66 92.92 17.15 22.95 13.57 79.16 90.90 89.78 86.73 14.02 17.00 12.42
DRAEM + SeaS 98.64 99.40 97.89 98.11 76.55 72.70 58.87 88.12 87.04 83.04 98.45 49.05 48.62 35.00 85.45 93.58 90.85 95.43 20.09 26.10 17.07
GLASS [9] 99.92 99.98 99.60 99.27 74.09 70.42 57.14 97.68 96.89 93.03 98.47 45.58 48.39 39.92 92.34 96.85 93.37 98.46 48.46 49.13 45.03
GLASS + SeaS 99.97 99.99 99.81 99.29 76.82 72.38 57.45 97.88 97.39 93.21 98.43 48.06 49.32 40.00 92.95 97.38 93.21 98.73 48.55 49.28 46.02
Average [41, 9] 98.96 99.22 97.97 98.59 70.99 68.23 58.72 91.98 91.10 87.35 95.70 31.37 35.67 26.75 85.75 93.88 91.58 92.60 31.24 33.07 28.73
Average(+ SeaS) 99.31 99.70 98.85 98.70 76.69 72.54 58.16 93.00 92.22 88.13 98.44 48.56 48.97 37.50 89.20 95.48 92.03 97.08 34.32 37.69 31.55
Table 3: Comparison on combining generated normal images with unsupervised anomaly detection methods across multiple datasets.
Segmentation Models MVTec AD VisA MVTec 3D AD
Image-level Pixel-level Image-level Pixel-level Image-level Pixel-level
AUROC AP $F_{1}$-max AUROC AP $F_{1}$-max IoU AUROC AP $F_{1}$-max AUROC AP $F_{1}$-max IoU AUROC AP $F_{1}$-max AUROC AP $F_{1}$-max IoU
HVQ-Trans [26] 96.38 98.09 95.30 97.60 47.95 53.32 45.03 90.11 88.18 84.08 98.10 28.67 35.05 24.03 68.15 84.38 85.20 96.40 17.23 24.59 20.51
HVQ-Trans + SeaS 97.25 98.48 95.78 97.58 48.53 53.84 44.61 92.12 90.35 86.23 98.15 29.52 36.00 23.60 71.26 90.35 89.23 96.56 19.34 26.40 20.47
PatchCore [32] 98.63 99.47 98.18 98.37 56.13 58.83 49.45 94.84 95.98 91.69 98.38 48.58 49.69 42.44 83.44 94.89 92.24 98.55 34.52 39.09 39.29
PatchCore + SeaS 98.64 99.48 98.22 98.37 63.98 64.07 55.43 94.97 96.06 91.81 98.41 48.60 49.72 42.46 83.88 94.97 92.32 98.56 34.65 39.41 39.43
MambaAD [15] 98.54 99.52 97.77 97.67 56.23 59.34 51.31 94.19 94.44 89.55 98.49 39.27 44.18 37.68 85.92 95.69 92.51 98.57 37.30 41.08 39.44
MambaAD + SeaS 98.80 99.64 98.40 97.66 56.86 59.70 51.51 94.23 94.65 89.93 98.70 39.33 43.99 36.62 88.67 96.60 93.41 98.74 35.46 39.59 39.51
Average [26, 32, 15] 97.85 99.03 97.08 97.88 53.44 57.16 48.60 93.05 92.87 88.44 98.32 38.84 42.97 34.72 79.17 91.65 89.98 97.84 29.68 34.92 33.08
Average(+ SeaS) 98.23 99.20 97.47 97.87 56.46 59.20 50.52 93.77 93.69 89.32 98.42 39.15 43.24 34.23 81.27 93.97 91.65 97.95 29.82 35.13 33.14
Table 4: Comparison on trained supervised segmentation models for anomaly detection and segmentation across multiple datasets.
Segmentation Models Generative Models MVTec AD VisA MVTec 3D AD
Image-level Pixel-level Image-level Pixel-level Image-level Pixel-level
AUROC AP $F_{1}$-max AUROC AP $F_{1}$-max IoU AUROC AP $F_{1}$-max AUROC AP $F_{1}$-max IoU AUROC AP $F_{1}$-max AUROC AP $F_{1}$-max IoU
BiSeNet V2 [40] DFMGAN 90.90 94.43 90.33 94.57 60.42 60.54 45.83 63.07 62.63 66.48 75.91 9.17 15.00 9.66 61.88 81.80 84.44 75.89 15.02 21.73 15.68
AnomalyDiffusion 90.08 94.84 91.84 96.27 64.50 62.27 42.89 76.11 77.74 73.13 89.29 34.16 37.93 15.93 61.49 81.35 85.36 92.39 15.15 20.09 14.70
SeaS 96.00 98.14 95.43 97.21 69.21 66.37 55.28 85.61 86.64 80.49 96.03 42.80 45.41 25.93 73.60 87.75 85.82 90.41 26.04 32.61 28.55
UPerNet [38] DFMGAN 90.74 94.43 90.37 92.33 57.01 56.91 46.64 71.69 71.64 70.70 75.09 12.42 18.52 15.47 67.56 84.53 84.99 75.12 19.54 26.04 18.78
AnomalyDiffusion 96.62 98.61 96.21 96.87 69.92 66.95 50.80 83.18 84.08 78.88 95.00 39.92 45.37 20.53 76.56 90.42 87.35 88.48 28.95 35.81 25.04
SeaS 98.29 99.20 97.34 97.87 74.42 70.70 61.24 90.34 90.73 84.33 97.01 55.46 55.99 35.91 82.57 92.59 88.72 91.93 38.51 43.53 38.56
LFD [45] DFMGAN 91.08 95.40 90.58 94.91 67.06 65.09 45.49 65.38 62.25 66.59 81.21 15.14 18.70 6.44 62.23 82.17 85.38 72.15 9.54 14.29 14.81
AnomalyDiffusion 95.15 97.78 94.66 96.30 69.77 66.99 45.77 81.97 82.36 77.35 88.00 30.86 38.56 16.61 77.06 89.44 87.20 92.68 24.29 32.74 19.90
SeaS 95.88 97.89 95.15 98.09 77.15 72.52 56.47 83.07 82.88 77.24 92.91 43.87 46.46 26.37 78.96 91.22 87.28 91.61 40.25 43.47 39.00
Average DFMGAN 90.91 94.75 90.43 93.94 61.50 60.85 45.99 66.71 65.51 67.92 77.40 12.24 17.41 10.52 63.89 83.83 84.94 74.39 14.70 20.69 16.42
AnomalyDiffusion 93.95 97.08 94.24 96.48 68.06 65.40 46.49 80.42 81.39 76.45 90.76 34.98 40.62 17.69 71.70 87.07 86.64 91.18 22.80 29.55 19.88
SeaS 96.72 98.41 95.97 97.72 73.59 69.86 57.66 86.34 86.75 80.69 95.32 47.38 49.29 29.40 78.38 90.52 87.27 91.32 34.93 39.87 35.37
@@
Evaluation Metrics. For image generation, unlike existing methods [11, 17] that only assess the whole anomaly images, our evaluation contains three levels: anomaly images, normal images, and anomalies, using 4 metrics: (1)Inception Score (IS) and Intra-cluster pairwise LPIPS distance (IC-LPIPS) [29] for authenticity and diversity of anomaly images. (2) KID [5] for authenticity of normal images. (3)IC-LPIPS calculated only in anomaly regions (short for IC-LPIPS(a)) for the diversity of anomalies. For pixel-level anomaly segmentation and image-level anomaly detection, we use 3 metrics: Area Under Receiver Operator Characteristic curve (AUROC), Average Precision (AP) and $F_{1}$-score at optimal threshold ($F_{1}$-max). We also report Intersection over Union (IoU) for segmentation.
@@ 4.2 Comparison in Anomaly Image Generation
Comparison Methods. For image generation, we compare SeaS with current anomaly image generation methods, like Crop&Paste [23], SDGAN [28], Defect-GAN [42](all without open-source code), DFMGAN [11], and AnomalyDiffusion [17] in terms of fidelity and diversity. For diverse generated anomalies, we combine SeaS-generated anomalies with synthesis-based AD approaches like DRAEM [41] and GLASS [9]. For authentic generated normal images, we use SeaS-generated normal images to augment the training sets of unsupervised AD methods like HVQ-Trans [26], PatchCore [32], and MambaAD [15]. For anomaly image-mask pairs, we generate them with DFMGAN, AnomalyDiffusion, and SeaS, to train segmentation models like BiSeNet V2 [40], UPerNet [38], and LFD [45] respectively. Different from AnomalyDiffusion, which trains one segmentation model per product, we train a unified supervised segmentation model for all products, which is more challenging.
@@
Anomaly image generation quality. In Tab. 1, we compare SeaS with some state-of-the-art anomaly image generation methods on fidelity (IS and KID) and diversity (IC-LPIPS and IC-LPIPS(a)). SeaS outperforms other methods in IS and IC-LPIPS, showing superior fidelity and diversity. It also excels in generating authentic normal images and diverse anomalies. Compared to AnomalyDiffusion, which cannot generate normal images, SeaS leads in IC-LPIPS(a). SeaS also surpasses DFMGAN in both KID and IC-LPIPS(a). We exhibit the generated anomaly images in Fig. 5, SeaS-generated anomaly images have higher fidelity (e.g., hazelnut_crack). Compared with other methods, SeaS can generate images with different types, colors, and shapes of anomalies rather than overfitting to the training images (e.g., foam_color). SeaS-generated masks are also precisely aligned with the anomaly regions (e.g., toothbrush_defective). We also present the authentic generated normal images in Fig. 6. More qualitative and quantitative anomaly image generation results are in appendix A.6.
Combining generated anomalies with synthesis-based AD methods. We replace the synthesized pseudo-anomalies in DRAEM [41] and GLASS [9] with SeaS-generated anomalies. As shown in Tab. 2, SeaS-generated anomalies, which offer sufficient diversity, consistently improve synthesis-based AD methods by suppressing false negatives, leading to better performance across multiple datasets. More training details are given in appendix A.3.
Combining generated normal images with AD methods. We use SeaS-generated normal images to supplement the training sets of existing state-of-the-art unsupervised AD methods, the results are given in Tab. 3. Using SeaS-generated normal images with minor local variations and global consistency, unsupervised AD methods reduce false positives and perform well across multiple metrics, improving industrial anomaly detection on various datasets. More training details are given in appendix A.3 .
@@
Training supervised segmentation models for anomaly segmentation and detection. We generate 1,000 image-mask pairs for each anomaly type and use them, along with all normal images in the original training sets, to train a unified supervised segmentation model. The models are tested on the remaining images not included in the training set. All methods are trained using the same number of images and the training settings, detailed in appendix A.4. As shown in Tab. 4, the segmentation results consistently demonstrate that our method outperforms others across all the segmentation models, with average IoU improvements of 11.17% (MVTec AD), 11.71% (VisA), and 15.49% (MVTec 3D AD). Segmentation anomaly maps are shown in Fig. 7. Using our generated image-mask pairs to train BiSeNet V2, there are fewer false positives in wood_combined and fewer false negatives in bottle_contamination. We also use the maximum value of the segmentation anomaly map as the image-level anomaly score for anomaly detection, achieving gains of 2.77% (MVTec AD), 5.92% (VisA), and 6.68% (MVTec 3D AD) in image-AUROC. More qualitative comparison results are in appendix A.7 and appendix A.8.
@@
Figure 7: Qualitative supervised anomaly segmentation results with BiSeNet V2 on MVTec AD.